# DvD: Unleashing a Generative Paradigm for Document Dewarping via Coordinates-based Diffusion Model

### v1

@@ Abstract.
Document dewarping aims to rectify deformations in photographic document images, thus improving text readability, which has attracted much attention and made great progress, but it is still challenging to preserve document structures. Given recent advances in diffusion models, it is natural for us to consider their potential applicability to document dewarping. However, it is far from straightforward to adopt diffusion models in document dewarping due to their unfaithful control on highly complex document images (e.g., 2000$\times$3000 resolution). In this paper, we propose DvD, the first generative model to tackle document Dewarping via a Diffusion framework. To be specific, DvD introduces a coordinate-level denoising instead of typical pixel-level denoising, generating a mapping for deformation rectification. In addition, we further propose a time-variant condition refinement mechanism to enhance the preservation of document structures. In experiments, we find that current document dewarping benchmarks can not evaluate dewarping models comprehensively. To this end, we present AnyPhotoDoc6300, a rigorously designed large-scale document dewarping benchmark comprising 6,300 real image pairs across three distinct domains, enabling fine-grained evaluation of dewarping models. Comprehensive experiments demonstrate that our proposed DvD can achieve state-of-the-art performance with acceptable computational efficiency on multiple metrics across various benchmarks including DocUNet, DIR300, and AnyPhotoDoc6300. The new benchmark and code will be publicly available.
@@
submissionid: aaaaccs: Applied computing Document analysisccs: Applied computing Optical character recognitionccs: Applied computing Document scanningccs: Computing methodologies Image processing
@@ 1. Introduction
Document dewarping has evolved through two distinct periods. The early model-based methods (Kanungo et al., 1993; Cao et al., 2003a; Liang et al., 2008) basically follow a ”reconstruct-then-dewarp” paradigm, which is typically limited by hardware dependence or surface assumptions. Subsequent data-driven methods (Ma et al., 2018; Li et al., 2019; Das et al., 2019) have shifted toward a mapping regression paradigm to model the condition probability. Specifically, given a large-scale photographic document dataset, these methods predominantly train neural networks to directly regress mapping (e.g., backward mapping, forward mapping) for dewarping. Accordingly, they also transfer the research attention to high-fidelity training data (Zhang et al., 2024; Verhoeven et al., 2023) and better network architecture for feature extraction (Li et al., 2023b; Feng et al., 2022).
@@
To make the evaluation of dewarping models fairly, we collect most of publicly available document dewarping benchmarks as shown in Table 1. We find that these benchmarks typically suffer from three critical limitations that impede a comprehensive evaluation of dewarping models, including restricted coverage of real-world scenarios, small scale of dataset size, and deficient annotation of domains. These limitations might have led to evaluation ambiguity, restricting the model’s application in real-world scenarios. To this end, we construct a fine-grained and large-scale benchmark AnyPhotoDoc6300, which contains 6,300 real-world photographic image document pairs, rigorously organized across three distinct domains, enabling fine-grained evaluation of dewarping models. In addition to the benchmark, we extend the evaluation metrics. We find that existing methods commonly utilize off-the-shelf OCR engines to measure the Edit Distance (ED) and Character Error Rate (CER) as OCR metrics (Das et al., 2019). We identify that there is still no exploration about whether the dewarped document can attain equivalent text readability to its flat counterpart for MLLMs. To fill the blank, we propose two MLLM-based OCR metrics (MMCER and MMED) in this work.
@@
  • Pioneering a paradigm shift, we present DvD, the first generative model to tackle document dewarping via a coordinates-based diffusion framework. Unlike typical pixel-level denoising to generate dewarped images directly, we operate coordinate-level denoising to generate coordinates mappings for dewarping.
@@ 2.1. Early Model-based Methods
Early model-based methods basically adhere to a two-step ”reconstruct-then-dewarp” paradigm. At the first step, for surface reconstruction, some methods utilize specialized hardware such as structured-light devices (Brown and Seales, 2001; Meng et al., 2014), laser scanners (Zhang et al., 2008), multi-view camera systems (Ulges et al., 2004; Koo et al., 2009; You et al., 2017a), and proximal light sources (Wada et al., 1997). Another methods bypass the hardware dependence by leveraging 2D visual cues (e.g., text lines (Tian and Narasimhan, 2011), local text orientation (Meng et al., 2018), text blocks (Kim et al., 2015), etc.) to estimate 3D geometry with parametric assumptions. In the second step, hand-crafted transformations for dewarping the surface to the plane are undertaken according to the reconstructed surface. For parameterized surfaces, typical transformations involve Generalized Cylinders Surface (GCS) (Cao et al., 2003b; Koo et al., 2009), developable surfaces (Liang et al., 2008), generalized ruled surfaces (Tsoi and Brown, 2007), and Non-Uniform Rational B-Splines (NURBS) (Tan et al., 2005). For non-parameterized surfaces, techniques such as planar strips (Meng et al., 2015), stiff mass-spring systems (Zhang et al., 2008), conformal mapping (You et al., 2017a), and sparse correspondence (Meng et al., 2018) are harnessed to tackle the mesh representation.
While effective in limited scenarios, these methods exhibited poor generalization to real-world warping patterns due to confined assumptions. Meanwhile, demanding hardware-dependent solutions also deviate from the prevalent habit of using mobile phones to digitize documents.
@@ 2.2. Data-driven Methods
Data-driven methods bring a shift toward a regression-based paradigm, significantly lessening reliance on both assumptions and hardware. Ma et al. (2018) pioneer to employ deep neural networks to solve document dewarping by framing it as a mapping regression task. Subsequent DewarpNet (Das et al., 2019) reformulates the traditional ”reconstruct-then-dewarp” paradigm via two regression networks while collecting a high-quality synthetic dataset Doc3D using rendering software. CREASE (Markovitz et al., 2020) explores multi-modal warped document representations to strengthen the dewarping performance. DocProj (Li et al., 2019) and PW-DewarpNet (Das et al., 2021) opt to dewarp documents in a patch-wise manner, benefiting the results at local details. DispFlow (Xie et al., 2020) and DDCP (Xie et al., 2021) investigate the effectiveness of mapping and sparse control points as deformation representations, respectively. DocTr (Feng et al., 2021) replaces convolutional network with vision transformer, achieving significant performance boosts. RDGR (Jiang et al., 2022), DocGeoNet (Feng et al., 2022), and FTA (Li et al., 2023b) leverage text-line features to keep textual content alignment. PaperEdge (Ma et al., 2022), Marior (Zhang et al., 2022), and DocReal (Yu et al., 2024) devise two-stage networks, which first coarsely dewarp the mask of document, then refine details. FDRNet (Xue et al., 2022) introduces frequency-domain insight, designing an image-level loss based on high-frequency textures extracted from the Fourier transformation. UVDoc (Verhoeven et al., 2023) exploits a novel data annotation pipeline through optical invisibility of ultraviolet ink to acquire pseudo-real training data. LA-Doc (Li et al., 2023a) constructs a synthetic dataset with large-scale deformation patterns via physical mass-spring systems and proposes a layout-aware dewarping network. Inv3D and DocMatcher (Hertlein et al., 2025, 2023) present a template-based document dewarping using auxiliary template information. These methods concentrate on curating training datasets and devising regression networks, whereas their intrinsic discriminative nature is unexplored, which constrains the dewarping performance.
@@ 3.2. Coordinates-Based Diffusion Model
Latent Diffusion Model in Coordinate Space. Directly generating $M_{0}$ explicit foster deformation-aware modeling to achieve better structural preservation, however both $M_{0}$ and $I_{0}$ actually contain equally high-resolution (e.g., 2000×3000), leading to high computational complexity. Inspired by the latent diffusion model (Rombach et al., 2021), we implement the diffusion and denoising processes only in a smaller coordinate space (i.e., $64\times 64$), significantly reducing the training and inference costs. Figure 2 shows the overall framework of our proposed DvD model. Concretely, guided by a compound condition $c_{t}$, our DvD progressively generates a series of latent variables $m$ from a random Gaussian distribution. This paradigm relies on a tailored conditional denoising diffusion probabilistic framework, which typically contains both forward and reverse processes (Ho et al., 2020; Song and Ermon, 2019). We define the forward diffusion process for mapping as the Gaussian transition, s.t. $q(m_{t}|m_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}m_{t-1},\beta_{t}\mathbf{I})$, where ${\beta_{t}}$ is a predefined variance schedule. The resulting latent variable $m_{t}$ can be formulated as:
@@
Compound Conditions $c_{t}$. Following previous works (Feng et al., 2022; Li et al., 2023b; Jiang et al., 2022), our DvD utilizes pre-trained multiple feature extractors (MFE) to enhance document visual perception and composes these features into a compound condition for the diffusion model, denoted as $c_{t}=\{f_{d},f_{m},f_{l},r_{t}\}$, where $f_{d},f_{m},f_{l}$ represent the features for raw document images, document foreground mask and text-lines, respectively. Note that $f_{d},f_{m},f_{l}$ are time-fixed conditions, while $r_{t}$ represents a time-variant condition that enables dynamic guidance in the denoising generation process.
Time-variant Condition Refinement (TVCR) Mechanism. As illustrated in the right region of Fig. 2, the intermediate dewarping results reveal the visual gap from intermediate denoising states to the ideal dewarped result. To harness this information for enhanced the preservation of document structures, we introduce a time-variant condition refinement mechanism within the reverse diffusion process to ensure increasingly precise document structure as the process evolves. Specifically, we incorporate a time-variant condition $r_{t}=\{m_{0|t},f_{0|t}\}$ for iteratively updating $r_{t}$, where $m_{0|t}$ means the predicted latent variables mapping in each time-step, and $f_{0|t}$ indicates dewarped document features $f_{d}$ using $m_{0|t}$ by Equ. 1. Unlike the time-fixed condition applied in vanilla diffusion models, the proposed time-variant condition reflects the intermediate variable $m_{0|t}$ and local structure deformation $f_{0|t}$ in each diffusion step, which facilitates the model to dynamically approach a tighter evidence lower bound (ELBO) (Rezende et al., 2014) for maximizing the conditional likelihood $\log p(m_{0}|\{c_{t}\})$.
Figure 4. A sample data array visualization across 3 distinct domain spectra (A, B, C) in our AnyPhotoDoc6300 Benchmark. We aim to evaluate dewarping capability of the model to different warping patterns (D) under well-distinguished domain combinations.
@@ 3.3. Network Architecture
In this section, we discuss the design of the network architecture in Fig. 2, including the multiple feature extractors (MFE) and diffusion decoder. In MFE, we employ three parallel networks to extract features $f_{d},f_{m},f_{l}$ for raw document image, document foreground, and text lines, respectively. For document image features $f_{d}$, we utilize the first three layers of VGG16 (Simonyan and Zisserman, 2014) to extract features. For document foreground features $f_{m}$, we concatenate the last six layers of U2Net (Qin et al., 2020) as a foreground segmentation network to extract those features associated with foreground. For text lines features $f_{l}$, we use the last decoder layer of the UNet (Ronneberger et al., 2015), which is pre-trained for text-line segmentation. All extracted features are uniformly downsampled to $64\times 64$ resolution compatible with the coordinate space. Furthermore, the diffusion decoder comprises 12 condition embedding blocks (CEB) and 6 fusion generation blocks (FGB). In each condition embedding block, we extend the architecture of diffusion transformers (DiT) (Peebles and Xie, 2023) to implement both time embeddings and cross-attention conditioning. The three features $f_{d},f_{m},f_{l}$ from the MFE and time-variant conditions ($r_{t}$) are jointly conditioned with the noisy latent variable $m_{t}$ through four parallel cross-attention computations. In the FGB, the results of the cross-attention computations are concatenated and fed into a six-layer FGB that contains self-attention and feed-forward layers. Subsequently, we apply three linear layers to obtain the denoised mapping $m_{0}$. Next, we upsample $m_{0}$ to obtain the high-resolution mapping $M_{0}$, which is used to produce the final flat document image $I_{0}$ via Equ. 1. In Appendix B, we provide more detailed architectures.
@@ 3.4. Training and Sampling
DvD Training Algorithm. Vanilla diffusion model randomly samples timesteps from a uniform distribution during the training phase (Ho et al., 2020), which mismatch the sequential acquisition of time-variant conditions. To solve this issue, we extend the vanilla diffusion model by integrating our TVCR mechanism, as detailed in Algorithm LABEL:alg:training, where we implement a certain number of sampling based on current timestep, thereby obtaining the intermediate dewarping latent variables $m_{0|t}$ and the corresponding intermediate dewarped features $f_{0|t}$.
Model Optimization. During the training phase, we freeze the pre-trained weights of the foreground and text-line segmentation network from DocGeoNet (Feng et al., 2022), including U2Net and UNet. However, for the VGG dedicated to document feature extraction raw document feature extraction, we jointly optimize it with the subsequent diffusion decoder. To optimize our DvD, instead of predicting noise in (Ho et al., 2020), we follow Luo et al. (2024) and Nam et al. (2024) to predict the generated object itself. Thus, the loss function is given by:
@@
Stochastic Sampling Property. As shown in Equ. 3, the reverse process of DDIM (Nichol and Dhariwal, 2021) introduces $\sigma$ to inject stochasticity into the sampling trajectory. To account for this property and enhance generation stability, we implement a dual-hypothesis strategy that simultaneously generates two candidate mappings. Afterward, we calculate the mean of the two mappings as a final result. We provide further training details in Appendix A.

### v2

@@ Abstract.
[…]
@@
submissionid: 1719journal: TOGccs: Applied computing Document analysisccs: Applied computing Optical character recognitionccs: Applied computing Document scanningccs: Computing methodologies Image processing
@@ 1. Introduction
Document dewarping has evolved through two distinct periods. The early model-based methods (Kanungo et al., 1993; Cao et al., 2003a; Liang et al., 2008) basically follow a ”reconstruct-then-dewarp” paradigm, which is typically limited by hardware dependence or surface assumptions. Subsequent data-driven methods (Ma et al., 2018; Li et al., 2019; Das et al., 2019) have shifted toward a mapping regression paradigm to model the condition probability. Specifically, given a large-scale photographic document dataset, these methods predominantly train neural networks to directly regress mapping (e.g., backward mapping, forward mapping) for dewarping. Accordingly, they also transfer the research attention to high-fidelity training data (Zhang et al., 2024c; Verhoeven et al., 2023; Zhang et al., 2024b) and better network architecture for feature extraction (Li et al., 2023b; Feng et al., 2022).
@@
To make the evaluation of dewarping models fairly, we collect most of publicly available document dewarping benchmarks as shown in Table 1. We find that these benchmarks typically suffer from three critical limitations that impede a comprehensive evaluation of dewarping models, including restricted coverage of real-world scenarios, a small dataset scale, and deficient annotation of domains. These limitations might have led to evaluation ambiguity, restricting the model’s application in real-world scenarios. To this end, we construct a fine-grained and large-scale benchmark AnyPhotoDoc6300, which contains 6,300 real-world photographic image document pairs, rigorously organized across three distinct domains, enabling fine-grained evaluation of dewarping models. In addition to the benchmark, we extend the evaluation metrics. We find that existing methods commonly utilize off-the-shelf OCR engines to measure the Edit Distance (ED) and Character Error Rate (CER) as OCR metrics (Das et al., 2019). We identify that there is still no exploration about whether the dewarped document can attain equivalent text readability to its flat counterpart for MLLMs. To fill the blank, we propose two MLLM-based OCR metrics (MMCER and MMED) in this work.
@@
  • Pioneering a paradigm shift, we present DvD, the first generative model to tackle document dewarping via a coordinates-based diffusion framework. Unlike typical pixel-level denoising to generate dewarped images directly, we operate coordinate-level denoising to generate coordinate mappings for dewarping.
@@ 2.1. Early Model-based Methods
Early model-based methods basically adhere to a two-step ”reconstruct-then-dewarp” paradigm. At the first step, for surface reconstruction, some methods utilize specialized hardware such as structured-light devices (Brown and Seales, 2001; Meng et al., 2014), laser scanners (Zhang et al., 2008), multi-view camera systems (Ulges et al., 2004; Koo et al., 2009; You et al., 2017a), and proximal light sources (Wada et al., 1997). Another methods bypass the hardware dependence by leveraging 2D visual cues (e.g., text lines (Tian and Narasimhan, 2011), local text orientation (Meng et al., 2018), text blocks (Kim et al., 2015), etc.) to estimate 3D geometry with parametric assumptions. In the second step, hand-crafted transformations for dewarping the surface to the plane are undertaken according to the reconstructed surface. For parameterized surfaces, typical transformations involve Generalized Cylinders Surface (GCS) (Cao et al., 2003b; Koo et al., 2009), developable surfaces (Liang et al., 2008), generalized ruled surfaces (Tsoi and Brown, 2007), and Non-Uniform Rational B-Splines (NURBS) (Tan et al., 2005). For non-parameterized surfaces, techniques such as planar strips (Meng et al., 2015), stiff mass-spring systems (Zhang et al., 2008), conformal mapping (You et al., 2017a), and sparse correspondence (Meng et al., 2018) are harnessed to tackle the mesh representation. While effective in limited scenarios, these methods exhibited poor generalization to real-world warping patterns due to confined assumptions. Meanwhile, demanding hardware-dependent solutions also deviates from the prevalent habit of using mobile phones to digitize documents.
@@ 2.2. Data-driven Methods
Data-driven methods bring a shift toward a regression-based paradigm, significantly lessening reliance on both assumptions and hardware. Ma et al. (2018) pioneer the application of deep neural networks to solve document dewarping by framing it as a mapping regression task. Subsequent DewarpNet (Das et al., 2019) reformulates the traditional ”reconstruct-then-dewarp” paradigm via two regression networks while collecting a high-quality synthetic dataset Doc3D using rendering software. CREASE (Markovitz et al., 2020) explores multi-modal warped document representations to strengthen the dewarping performance. DocProj (Li et al., 2019) and PW-DewarpNet (Das et al., 2021) opt to dewarp documents in a patch-wise manner, benefiting the results at local details. DispFlow (Xie et al., 2020) and DDCP (Xie et al., 2021) investigate the effectiveness of mapping and sparse control points as deformation representations, respectively. DocTr (Feng et al., 2021) replaces the convolutional network with a vision transformer, achieving significant performance boosts. RDGR (Jiang et al., 2022), DocGeoNet (Feng et al., 2022), and FTA (Li et al., 2023b) leverage text-line features to keep textual content alignment. PaperEdge (Ma et al., 2022), Marior (Zhang et al., 2022), and DocReal (Yu et al., 2024) devise two-stage networks, which first coarsely dewarp the mask of document, then refine details. FDRNet (Xue et al., 2022) introduces frequency-domain insight, designing an image-level loss based on high-frequency textures extracted from the Fourier transformation. UVDoc (Verhoeven et al., 2023) exploits a novel data annotation pipeline through optical invisibility of ultraviolet ink to acquire pseudo-real training data. LA-Doc (Li et al., 2023a) constructs a synthetic dataset with large-scale deformation patterns via physical mass-spring systems and proposes a layout-aware dewarping network. Inv3D and DocMatcher (Hertlein et al., 2025, 2023) present a template-based document dewarping using auxiliary template information. DocHFormer (Zhou et al., 2025) introduces a novel shuffle transformer block to harmonize feature representation. DocRes (Zhang et al., 2024a) develops a unified model that integrates multiple low-quality document enhancement tasks. These methods concentrate on curating training datasets and devising regression networks, whereas their intrinsic discriminative nature is unexplored, which constrains the dewarping performance.
@@ 3.2. Coordinates-Based Diffusion Model
3.2.1. Latent Diffusion Model in Coordinate Space.
Directly generating $M_{0}$ explicit foster deformation-aware modeling to achieve better structural preservation, however both $M_{0}$ and $I_{0}$ actually contain equally high-resolution (e.g., 2000×3000), leading to high computational complexity. Inspired by the latent diffusion model (Rombach et al., 2021), we implement the diffusion and denoising processes only in a smaller coordinate space (i.e., $64\times 64$), significantly reducing the training and inference costs. Figure 2 shows the overall framework of our proposed DvD model. Concretely, guided by a compound condition $c_{t}$, our DvD progressively generates a series of latent variables $m$ from a random Gaussian distribution. This paradigm relies on a tailored conditional denoising diffusion probabilistic framework, which typically contains both forward and reverse processes (Ho et al., 2020; Song and Ermon, 2019). We define the forward diffusion process for mapping as the Gaussian transition, s.t. $q(m_{t}|m_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}m_{t-1},\beta_{t}\mathbf{I})$, where $\beta_{t}$} is a predefined variance schedule. The resulting latent variable $m_{t}$ can be formulated as:
@@
Figure 4. Detailed architectures of condition embedding blocks (CEB) and fusion generation blocks (FGB). We ignore some simple operations, such as cosine position encoding and activation functions.
3.2.2. Compound Conditions $c_{t}$.
Following previous works (Feng et al., 2022; Li et al., 2023b; Jiang et al., 2022), our DvD utilizes pre-trained multiple feature extractors (MFE) to enhance document visual perception and composes these features into a compound condition for the diffusion model, denoted as $c_{t}=\{f_{d},f_{m},f_{l},r_{t}\}$, where $f_{d},f_{m},f_{l}$ represent the features for raw document images, document foreground mask and text-lines, respectively. Note that $f_{d},f_{m},f_{l}$ are time-fixed conditions, while $r_{t}$ represents a time-variant condition that enables dynamic guidance in the denoising generation process.
3.2.3. Time-variant Condition Refinement (TVCR) Mechanism.
As illustrated in the right region of Fig. 2, the intermediate dewarping results reveal the visual gap from intermediate denoising states to the ideal dewarped result. To harness this information for enhanced preservation of document structures, we introduce a time-variant condition refinement mechanism within the reverse diffusion process to ensure increasingly precise document structure as the process evolves. Specifically, we incorporate a time-variant condition $r_{t}=\{m_{0|t},f_{0|t}\}$ for iteratively updating $r_{t}$, where $m_{0|t}$ means the predicted latent variables mapping in each time-step, and $f_{0|t}$ indicates dewarped document features $f_{d}$ using $m_{0|t}$ by Equ. 1. Unlike the time-fixed condition applied in vanilla diffusion models, the proposed time-variant condition reflects the intermediate variable $m_{0|t}$ and local structure deformation $f_{0|t}$ in each time-step, which facilitates the model to dynamically approach a tighter evidence lower bound (ELBO) (Rezende et al., 2014) for maximizing the conditional likelihood $\log p(m_{0}|\{c_{t}\})$.
Figure 5. A sample data array visualization across 3 distinct domain spectra (A, B, C) in our AnyPhotoDoc6300 Benchmark. We aim to evaluate the dewarping capability of the model to different warping patterns (D) under well-distinguished domain combinations.
@@ 3.3. Network Architecture
In this section, we discuss the design of the network architecture in Fig. 2, including the multiple feature extractors (MFE) and diffusion decoder. In MFE, we employ three parallel networks to extract features $f_{d},f_{m},f_{l}$ for raw document image, document foreground, and text lines, respectively. For document image features $f_{d}$, we utilize the first three blocks of VGG16 (Simonyan and Zisserman, 2014) to extract features. For document foreground features $f_{m}$, we concatenate the last six layers of U2Net (Qin et al., 2020) as a foreground segmentation network to extract those features associated with foreground. For text lines features $f_{l}$, we use the last decoder layer of the UNet (Ronneberger et al., 2015), which is pre-trained for text-line segmentation. All extracted features are uniformly downsampled to $64\times 64$ resolution compatible with the coordinate space. Furthermore, as shown in Fig. 4, the diffusion decoder comprises 12 condition embedding blocks (CEB) and 6 fusion generation blocks (FGB). In each CEB, we extend the standard architecture of diffusion transformers (DiT), i.e. DiT_S_2 (Peebles and Xie, 2023) to implement both time embeddings and cross-attention conditioning. To embed input time-steps, we adopt the same two-layer MLP in standard DiT to represent a 256-dimensional frequency embedding. To enable the condition control, we extend a multi-head cross-attention mechanism, where the compound condition $c_{t}$ serves as the key and value, and the noisy latent variable $m_{t}$ is used as the query. Specifically, we employ four parallel CEBs to decouple different conditions, including three feature conditions($f_{d},f_{m},f_{l}$) from the MFE and one time-variant condition ($r_{t}$). Subsequently, the CEB produces the hidden representations for the four conditions: $x_{d}$, $x_{m}$, $x_{l}$, and $x_{r}$. In the FGB, the results of CEB are concatenated and fed into the FGB, which contains self-attention and feed-forward layers. Subsequently, we apply three linear layers as a projector to obtain the denoised mapping $m_{0}$. Next, we directly upsample $m_{0}$ to obtain the high-resolution mapping $M_{0}$, which is used to produce the final flat document image $I_{0}$ via Equ. 1. In Appendix B, we provide more detailed architectures.
1:repeat
[…]
3:if $t=T$ then
4:  $r_{T}=\{O,O\}$, where $O$ represent all zeros for initialization
5:else if $t<T$ then
6:  for $t\in[T-1,t]$ do
7:   Sampling intermediate latent variable $m_{0|t}$ by Equ. 3
8:   Using $m_{0|t}$ to obtain intermediate dewarped
feature $f_{0|t}$ by Equ. 1
9:  end for
10:  $r_{t}=\{m_{0|t},f_{0|t}\}$
11:end if
[…]
13: Optimize $m_{0|t}=\mathcal{\epsilon}_{\theta}(m_{t},t,c_{t})$ by Equ. 4
14:until Converged
ALGORITHM 1 DvD training for TVCR mechanism
@@ 3.4. Training and Sampling
3.4.1. DvD Training Algorithm.
Vanilla diffusion model randomly samples time-steps from a uniform distribution during the training phase (Ho et al., 2020), which mismatches the sequential acquisition of time-variant conditions. To solve this issue, we extend the vanilla diffusion model by integrating our TVCR mechanism, as detailed in Algorithm 1, where we implement a certain number of sampling based on current time-step, thereby obtaining the intermediate dewarping latent variables $m_{0|t}$ and the corresponding intermediate dewarped features $f_{0|t}$.
3.4.2. Model Optimization.
During the training phase, we freeze the pre-trained weights of the foreground and text-line segmentation network from DocGeoNet (Feng et al., 2022), including U2Net and UNet. However, for the VGG dedicated to extracting raw document features, we jointly optimize it with the subsequent diffusion decoder. To optimize our DvD, instead of predicting noise in (Ho et al., 2020), we follow Luo et al. (2024) and Nam et al. (2024) to predict the generated object itself. Thus, the loss function is given by:
@@
3.4.3. Stochastic Sampling Property.
As shown in Equ. 3, the reverse process of DDIM (Nichol and Dhariwal, 2021) introduces $\sigma$ to inject stochasticity into the sampling trajectory. To account for this property and enhance generation stability, we implement a dual-hypothesis strategy that simultaneously generates two mappings. Afterward, we calculate the mean of the two mappings as a final result. We provide further training settings in Appendix A.